# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v1

@@ 3.2 Analysis of Non-Adaptive Compositions with Secure Summation
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2). Denote $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $H_{T}^{+}$ denotes the Moore–Penrose pseudoinverse of the matrix $H_{T}$. Then, from the point-of-view of a node $i$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
3.3 Analysis of Adaptive Compositions with Secure Summation
When analyzing adaptive compositions, the gossip averaging algorithm can be written as
$\theta_{t+1}=W(\theta_{t}+x_{t}(\theta_{t})+u_{t}),$ (3.3)
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $\theta_{t}$ denote the global state of the average estimate at round $t$. Naturally, each element of $x_{t}$ depends only on the corresponding element of $\theta_{t}$, i.e., for any $i\in[n]$, $[x_{t}(\theta_{t})]_{i}=[x_{t}([\theta_{t}]_{i})]_{i}$.
Central to our analysis is the observation that in case $D\simeq_{j}D^{\prime}$, in case of both $D$ and $D^{\prime}$, the view of each node $i$ in the graph is post-processing of the view of $\bar{N_{j}}$, the closed neighborhood of the node $j$, i.e., for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained from post-processing of the views $\mathrm{View}_{\mathcal{M}(D)}\big{(}\bar{N_{j}}\big{)}$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}\big{(}\bar{N_{j}}\big{)}$, respectively. As we show, this also allows analyzing adaptive compositions.
Notice that the selector matrix $S(\bar{N}_{j})$ selects the rows of $W$ corresponding to the indices in the set of closed neighborhood $\bar{N}_{j}$. More specifically, if $\left|\bar{N}_{j}\right|=d_{j}$ and $\bar{N}_{j}=\{i_{1},\ldots,i_{d_{j}}\}$, then
$S(\bar{N}_{j})W=\begin{bmatrix}e_{i_{1}}^{\top}W\\ \vdots\\ e_{i_{d_{j}}}^{\top}W\end{bmatrix}\in\mathbb{R}^{d_{j}\times n}.$
Denoting $S_{j}:=S(\bar{N}_{j})$, from the point-of-view of $\bar{N}_{j}$, what is observed is then given by $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$ (3.4)
$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
$\vdots$
$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})\cdots+W(x_{T-1}+u_{T-1})\big{)}.$
With a similar derivation as in the case of non-adaptive compositions, this representation gives the following result.
@@ 3.3 Analysis of Adaptive Compositions with Secure Summation
Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$. Denote
$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
We now turn to the analysis of node-level differential privacy guarantees in the absence of secure summation, where nodes can observe the plain messages sent by their neighbors. In this setting, we analyze both non-adaptive and adaptive compositions simultaneously. However, we note that for the non-adaptive case, our techniques can yield tighter bounds by incorporating the distance between the observing node and the node whose privacy is being evaluated.
3.4 Gossip Averaging without Secure Summation
In case the nodes are allowed to see the plain locally perturbed messages of their neighboring nodes, the analysis changes. Now, crucial observation is that the view of any node in the graph is post-processing of the view of the differing node only. I.e., in case $D\simeq_{j}D^{\prime}$, then for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained by applying the same $j$th node-independent post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{j\})$, respectively.
To analyze the view $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$, instead of the iteration (3.3), we have
$\mathrm{View}_{\mathcal{M}(D)}(\{j\})=\begin{bmatrix}e_{j}^{\top}\theta_{1}\\ \vdots\\ e_{j}^{\top}\theta_{T}\end{bmatrix}$
where
$\theta_{t+1}=W\theta_{t}+x_{t}+u_{t},\quad\theta_{0}=0,$ (3.5)
Writing out the recursion as in Eq. (3.4) and following the same steps, gives the following result.
Theorem 4.
Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
3.5 Accounting for Nodes’ Knowledge of Injected Noise
As stated in Section 2, we have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes. In practice, this likely has a small effect on the the DP guarantees. However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees.
For example, in case of non-adaptive compositions and secure summation, instead of Thm. 2, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big{(}T\cdot(n-1)\big{)}$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. Similar correction can be carried out also in all the other cases. We do this correction in all our experiments.
@@ 4 Experiments
Illustration of the DP Bounds for Non-Adaptive Compositions. We first consider numerical evaluation of the DP bounds applicable for non-adaptive compositions (Thm. 2 and Thm. 4).
First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$. We take the gossip matrix $W$ to be a doubly stochastic matrix obtained via a max degree normalization, where we first scale the off-diagonal elements as $W_{ij}=A_{ij}/\max\{d_{i},d_{j}\}$, where $A$ denotes the symmetric adjacency matrix and $d_{i}$ the degree of node $i$, $i\in[n]$. We then set the diagonal elements as $W_{ii}=1-\sum_{j,i\neq j}W_{ij}$. The left figure of Fig. 1 illustrates that in case of the secure summation, the squared sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}$ for a randomly chosen $j$ grows linearly w.r.t. $T$, and in particular, approaches $\frac{T}{n}$ which corresponds to the sensitivity in the case of centralized aggregation.
Second, consider the Facebook Social Circle dataset (Leskovec and Mcauley,, 2012) available in the Stanford Large Network Dataset Collection (Leskovec and Sosič,, 2016), which describes an directed graph of $n=4036$ nodes connected via binary edge weights. We consider an undirected version of this graph that we obtain via symmetrization of the adjacency matrix. We obtain a doubly stochastic gossip matrix $W$ using the max degree normalization similarly to the synthetic case. The right figure of Fig. 1 illustrates the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ for observing the node $j$ from node $i$, where $j$ is the node index 300 (an arbitrary node) and $i$ is either one of the first two indices in the neighborhood of node $j$ or one of the first two indices outside of the neighborhood of node $j$.
@@
Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: Facebook Social Circle dataset (Leskovec and Mcauley,, 2012), and the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ as a function of $T$. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not. In case of the synthetic Erdős–Rényi graph, the scaled sensitivity approaches $\frac{1}{n}$ which corresponds to the scaled sensitivity in the centralized case using secure aggregation.
Illustration of the DP Bounds for Adaptive Compositions. We also consider comparing the bounds in case of adaptive compositions (Thm. 3 and Thm. 4), which corresponds to gossip averaging based ML model training via private gradient descent, for example. We consciously exclude the composition bound of (Cyffers et al.,, 2022, Thm. 6) in our comparisons. Their bound exhibits an $O(T^{2})$ dependency in the Rényi DP parameters and becomes even looser than the standard LDP bound in our setting (when the number of Muffliato averaging steps in their algorithm is $K=1$). We compute average the $\varepsilon$-values of the pair-wise DP guarantees, i.e., the average of the DP-$\varepsilon$’s between all pairs $(i,j)$ over the whole graph, when $\delta=10^{-5}$. Interestingly, the bound we obtain using Thm. 4 for the averaging without secure summation becomes tighter than the bound for the averaging with secure summation (Thm. 3) as $T$ grows. This suggests that the bound of Thm. 3 can be improved further. Notably, both bounds represent clear improvements over the LDP bound.
Figure 2: DP amplification for adaptive compositions over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=50$ and $p=0.2$.
Decentralized Learning for Logistic Regression. We distribute the MNIST (LeCun,, 1998) dataset IID to 100 nodes and each node is minimizing the cross-entropy loss for the logistic regression model using gradient descent. The nodes are connected based on an adjacency matrix $A\in\{0,1\}^{n\times n}$ that corresponds to a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$. We compute the average DP-$\varepsilon$ values of all pairs $(i,j)$ (DP of node $j$ from the point of view of node $i$) over the whole graph, when $\delta=10^{-5}$. The messages of the nodes are locally DP updated models and nodes carry out gossip averaging over the closed neighborhoods, i.e., the row-stochastic gossip matrix $W=\Lambda^{-1}(A+I_{n})$, where $\Lambda$ is a diagonal matrix with $\Lambda_{ii}=\left|\bar{N_{i}}\right|^{-1}$, $i\in[n]$.
Figure 3: Decentralized learning for logistic regression on MNIST dataset over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$. The gossip matrix is the row-stochastic matrix corresponding to averaging over the closed neighborhood of each node.
5 Conclusions and Outlook
In this work, we have given the first privacy amplification results for DP gossip averaging algorithms on graphs that are able to accurately capture the privacy amplification arising from all the noise injected in the system. As a future work, it will be interesting to see, whether the strong results for non-adaptive compositions that use secure summation can be translated to bounds for adaptive compositions. From computational perspective, a necessary future task will be to speed up linear algebraic subroutines for evaluating the total sensitivities for large and possibly sparse graphs. Also, the proposed framework is likely applicable to other decentralized optimization methods exhibiting a similar linear structure as the gossip averaging considered here.

### v2

@@ 3.2 Analysis of Non-Adaptive Compositions
where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector. Then, from the point-of-view of a node $i$, the node $j$ is tightly $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
3.3 Accounting for Nodes’ Knowledge of Injected Noise
We have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes. In practice, this likely has a small effect on the the DP guarantees. However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees. Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by allouahprivacy is scheme called SecLDP.
For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm. 8, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. Similar correction can be carried out also in all the other cases. We do this correction in all our experiments.
3.4 Adaptive Compositions and Connection to Matrix-Mechanism Accounting
The recent work by bellet2025unified shows that the linear systems perspective given in Theorem 8 can be equivalently analyzed via a so-called matrix mechanisms (kairouz2021practical; pillutla2025correlated). This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis. We next shortly discuss this connection. At a high level, both frameworks reduce an observer’s view to a linear image of a data-dependent signal plus correlated Gaussian noise.
Different Threat Scenarios.
The matrix mechanism formulation by bellet2025unified allows analyzing private decentralized optimization methods that have a linear structure. Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
  • The LDP approach where the adversary sees all messages (bellet2018personalized). This corresponds to the setting of Theorem 8 with $S=I_{n}$.
  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (networkdp2). This corresponds to Theorem 8 with a selector matrix of a finite collection of nodes.
  • SecLDP (allouahprivacy), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees. This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Thmeorem 8, see Section 3.3.
Privacy Analysis of Matrix Mechanisms.
The matrix-mechanism analysis by bellet2025unified provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings. The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
In the matrix mechanism formulation, the $d$-dimensional user-wise messages are stacked into a global matrix $G\in\mathbb{R}^{nT\times d}$ of the form
$G=\begin{bmatrix}G_{1}\\ \vdots\\ G_{T}\end{bmatrix}$,
where each block $G_{t}\in\mathbb{R}^{n\times d}$ corresponds to round $t$ and is given by
$G_{t}=\begin{bmatrix}g_{t}^{1}\\ \vdots\\ g_{t}^{n}\end{bmatrix}$,
with $g_{t}^{i}$ representing the message of node $i$ at iteration $t$. Thus, differences between neighboring datasets appear as differences along the rows $(g_{t}^{i})_{t=1}^{T}$ corresponding to a single user.
The privacy analyses of different private optimization methods under varying threat scenarios are then analyzed via the following results.
Theorem 9 (Bellet et al., 2025, Thm. 5).
For each of the three trust models, LDP, PNDP and SecLDP, and for every linear decentralized learning algorithm, there exist matrices $A$, $B$ and $C$ such that the adversarial view can be written as
$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (3.3)
with $A=BC$.
To quantify privacy, one needs to bound how much $G$ may vary under the adjacency relation while remaining observable through this linear mapping.
Definition 3.1 (Sensitivity under participation).
Let $B$ and $C$ be two matrices, and let $\Pi$ be a participation scheme for $G$. The sensitivity of $C$ with respect to $B$ under $\Pi$ is defined as
$\mathrm{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\,\|\,C(G-G^{\prime})\,\|_{B^{+}B},$
where $\|v\|_{B^{+}B}:=\sqrt{v^{\top}(B^{+}B)v}$ denotes the seminorm induced by $B^{+}B$, and $G\simeq_{\Pi}G^{\prime}$ indicates that $G$ and $G^{\prime}$ differ only along entries that belong to the same individual according to $\Pi$.
A participation scheme $\Pi$ specifies which coordinates of the stacked update matrix $G$ may change under a neighboring-dataset relation. In other words, $\Pi$ encodes which entries of $G$ correspond to contributions of the same individual across time, ensuring that sensitivity is computed only over data differences permitted by the privacy definition.
Theorem 10 (Bellet et al., 2025, Thm. 7).
Let $\mathcal{O}_{\mathcal{A}}=AG+BZ$ be the attacker knowledge of a trust model on a linear decentralized learning algorithm, and denote by $\mathcal{M}(G)$ the corresponding mechanism. Let $\Pi$ be a participation scheme for $G$, and suppose $Z\sim\mathcal{N}(0,\nu^{2}I)$. Assume that $A$ is a column-echelon matrix and that $A=BC$ for some $C$. Then, if
$\nu=\sigma\,\mathrm{sens}_{\Pi}(C;B)\qquad\text{with}\qquad\mathrm{sens}_{\Pi}(C;B)\;\leq\;\max_{\pi\in\Pi}\;\sum_{s,t\in\pi}\bigl|\bigl(C^{\top}B^{+}BC\bigr)_{s,t}\bigr|,$
the mechanism $\mathcal{M}$ satisfies $\tfrac{1}{\sigma}$-Gaussian Differential Privacy, even when $G$ is chosen adaptively.
Adaptive Analysis.
Notably, bellet2025unified extend the existing adaptive analysis of matrix mechanisms (denisov2022improved) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns. The crucial implication is that once the representation
$\mathcal{O}_{\mathcal{A}}=AG+BZ$
is identified, privacy guarantees for adaptive compositions are the same as in the case of non-adaptive compositions and directly follow from a single sensitivity bound on $(B,C)$ under the participation scheme $\Pi$.
Matrix Mechanism Formulation of Gossip Averaging.
We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step. As shown in (Appendix A.4 bellet2025unified), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm. 8.
Equivalence of Sensitivity Bounds.
We also see from Thm. 10 that the sensitivity upper bound equals the upper bound given in our Thm. 8 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
$\left|C^{\top}B^{+}BC\right|=\left|VV^{\top}\right|$
and the summation in the upper bound of Thm. 10 equals the upper bound given in our Thm. 8 in case the differing node participates in every round.
4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
We first need the following auxiliary lemma.
Lemma 11.
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic. Then there exist constants $M<\infty$ and $\rho\in(0,1)$ such that for all $t\geq 1$
$W^{t}\;=\;\Pi\;+\;R^{t},\qquad\Pi:=\mathbf{1}\,\pi^{\top},\qquad\|R^{t}\|\;\leq\;M\,\rho^{t},$
where $\pi^{\top}W=\pi^{\top}$, $\pi>0$, $\sum_{i}\pi_{i}=1$.
Remark 12.
A row-stochastic primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value < 1, and as the proof of Lemma 11 shows, $1-\rho$ equals the spectral gap $\lambda_{1}-\max_{i\geq 2}\left|\lambda_{i}\right|$.
@@ 4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic with stationary distribution $\pi>0$, $\sum_{k=1}^{n}\pi_{k}=1$, $\pi^{\top}W=\pi^{\top}$. Let $H_{T}$ be the system matrix given in Lemma 7 with $S=e_{i}^{\top}$ (view of a single node, i.e., that of secure summation protocol) and define
$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
as in Theorem 8. Then
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
Remark 14.
Notice that if $W$ is doubly stochastic, $\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}=\frac{1}{n}$ and we asymptotically have the same sensitivity as in case of a central aggregator.
@@ 5 Experiments
Illustration of the DP Bounds for Secure Summation Protocols. We first consider numerical evaluation of the DP bounds we obtain from Thm. 8.
First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$. We take the gossip matrix $W$ to be a row-stochastic matrix obtained from row-wise averaging. Second, we consider a network generated using a preferential-attachment model (barabasi1999emergence) with $n=200$ nodes, initialized from a fully connected core of 5 nodes. Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. The gossip matrix $W$ is obtained from row-wise averaging.
As Figure 1 illustrates, in case of the secure summation, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ for a randomly chosen $j$ grows asymptotically linearly w.r.t. $T$ as predicted by Thm. 13.
@@
Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: gossip matrix of a preferential-attachment model with $n=200$ nodes, and the scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
6 Conclusions
We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that secure-summation-based averaging has asymptotic sensitivity comparable to central aggregation. Future work includes accelerating the linear-algebra routines needed to compute total sensitivities on large or sparse graphs and analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees.